# arXiv:2603.00175v1[cs.CV] 26 Feb 2026

### v1

@@ 3.3 Absorbing Markov Chain Interpretation of InfSA
The Neumann-series formulation of Pure InfSA (Eq. 10) admits a direct probabilistic interpretation through absorbing Markov chains [46, 31, 52]. This connection shows that InfSA computes expected visit counts in a random walk over the token graph, linking self-attention to classical stochastic processes.
@@
where $\mathbf{M}_{ij}\geq 0$ represents the transition probability from token $i$ to token $j$, and $R_{i}\geq 0$ is the per-step absorption probability at token $i$. The matrix $\mathbf{M}$ is substochastic ($\mathbf{M}\mathbf{1}\leq\mathbf{1}$, with strict inequality for at least one row) whenever $\gamma\max_{i}\sigma_{i}<1$, where $\sigma_{i}=\sum_{j}\hat{A}_{ij}$. Frobenius normalization constrains $\|\hat{A}\|_{F}=1$, which ensures $\rho(\mathbf{M})=\gamma\,\rho(\hat{A})<1$. We augment the $N$ token states with a single absorbing state $\mathfrak{a}$, yielding the canonical form [31]:
@@
which is precisely the Neumann kernel already underlying InfSA (Eq. 10). The entry $N_{ij}$ has a concrete probabilistic meaning [31, 52]: it equals the expected number of times the random walk visits token $j$ before absorption, given that it starts at token $i$. The InfSA path-integral matrix $S=N-I=(I-\gamma\,\hat{A})^{-1}-I$ therefore counts expected visits excluding the starting state.
@@
These scores coincide with Katz centrality [30], confirming that InfSA ranks tokens by global structural role rather than by local query–key affinity.
Why Frobenius normalization enables absorption. Softmax normalization gives row-stochastic matrices ($\hat{A}\mathbf{1}=\mathbf{1}$), corresponding to a closed Markov chain with no absorbing state—the source of oversmoothing [34]. Frobenius normalization breaks row-stochasticity ($\rho(\hat{A})<1$ in practice), introducing a positive absorption probability at every step and ensuring convergence. This absorbing-chain view is structurally identical to the Markov interpretation of Inf-FS [48, 46], where features are transient states and the fundamental matrix ranks them by expected walk persistence. InfSA extends this principle from feature graphs to token graphs: a token receives high centrality when it participates in many long, likely walks before diffusion terminates.
@@ 3.4 Linear-InfSA: Efficient Centrality Approximation
providing a positive signal that reflects token prominence in embedding space [50]. These energies are normalized without softmax:
@@
yielding a soft importance score that serves as a proxy for the dominant eigenvector of a positive operator [30, 26]. No softmax or Frobenius normalization is required; the $\ell_{1}$ constraint ensures numerical stability.
@@
Query-independent weighting by design. The weight vector $a$ depends only on Keys (equivalently Queries, since Q=KQ{=}K) and is shared across all positions—a direct consequence of the eigenvector limit, where the operator is dominated by its principal eigenvector $v$, producing an effective rank-1 matrix $vv^{\top}$. We recover expressivity by using 64 heads (vs. 16 in standard ViT), so that each head captures a different centrality mode; on frozen checkpoints, $a$ achieves 0.985 mean cosine similarity with the true Perron eigenvector (Sec. 4.5). Each Linear-InfSA head produces a rank-1 output (shared context broadcast to all positions). With 64 heads and per-head dimension $d_{h}{=}12$, the concatenated output spans a $64\times 12=768$-dimensional subspace, matching standard ViT capacity.
Linear-InfSA avoids constructing the full attention matrix, reducing complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. The vector $a$ is a linear-time surrogate for the dominant eigenvector [52, 33], obtained classically as $v=\lim_{k\to\infty}A^{k}x_{0}/\|A^{k}x_{0}\|_{1}$; this $k{\to}\infty$ characterization motivates the term “infinite” in Linear-InfSA. The ReLU in Eq. (19) suppresses anti-aligned tokens, and under nonnegativity the map defines a positively 1-homogeneous operator whose iterates converge to a unique direction by nonlinear Perron–Frobenius theory [33, 4] (see Sec. 3.6 for the full convergence argument).
@@
Relation to global pooling. Although the per-head output is a broadcast context vector, Linear-InfSA is not reducible to generic global pooling [27] or additive-attention aggregation [65]: the weight vector $a$ is the unique fixed point of a nonlinear Perron–Frobenius operator whose spectrum encodes the full inter-token similarity structure, not a learned or heuristic saliency score. This spectral grounding is verified empirically—$a$ achieves 0.985 mean cosine similarity with the true Perron eigenvector (Sec. 4.5)—and ensures that the aggregation captures global graph centrality rather than local token prominence.
3.5 Why Pure InfSA and Linear-InfSA Are “Infinite”
@@
3.6 Convergence of Linear-InfSA
@@
By the nonlinear Perron–Frobenius theorem for positively 1-homogeneous, order-preserving maps on the positive cone [33, 4], the normalized iterates $F^{k}(x_{0})/\|F^{k}(x_{0})\|_{1}$ converge to a unique eigenvector direction. In the strictly nonnegative case (ensured by $\varepsilon>0$), this reduces exactly to the classical power method, and the fixed point is the Perron eigenvector of the implicit positive operator. In all experiments we use $\varepsilon=10^{-6}$. Our empirical validation (Sect. 4.5) confirms that the single-step Linear-InfSA weight vector $a$ achieves 0.985 mean cosine similarity with this eigenvector.
4 Experiments
We evaluate InfSA by integrating it into a 4-layer Vision Transformer (ViT) with patch size 16. Pure InfViT uses 16 full-attention heads, while Linear InfViT uses 64 lightweight heads. Both models have comparable compute and parameter counts (58M vs. 53M). Full training details are provided in Sec. 4.7.
4.1 Scalability to Extreme Input Resolutions
[…]
Train Infer Train Infer Train Infer
24-layer, 16-head configuration ($d_{h}{=}48$) — fair depth comparison
Linformer [61] L $\mathcal{O}(Nd)$ 331.5M 55.64 38.95 17.97 25.67 16.69 7.79 10242
Performer [10] L $\mathcal{O}(Nd^{2})$ 330.3M 58.14 42.80 17.20 23.36 17.44 8.56 10242
SOFT [38] L $\mathcal{O}(Nd)$ 330.8M 56.10 36.44 17.83 27.44 16.83 7.29 10242
FLatten [21] L $\mathcal{O}(Nd^{2})$ 330.5M 57.82 40.15 17.29 24.91 17.35 8.03 10242
Agent Attn [23] L $\mathcal{O}(Nd)$ 331.0M 56.55 35.81 17.68 27.93 16.97 7.16 10242
MLLA [22] L $\mathcal{O}(Nd^{2})$ 330.4M 57.38 39.52 17.43 25.30 17.21 7.90 10242
InfViT Linear I $\bm{\mathcal{O}(N)}$ 305.4M 55.52 25.28 18.01 39.56 16.66 5.06 92162
4-layer, 64-head configuration ($d_{h}{=}12$) — proposed lightweight
Linformer [61] L $\mathcal{O}(Nd)$ 58.3M 12.49 5.63 80.04 177.62 3.75 1.13 10242
Performer [10] L $\mathcal{O}(Nd^{2})$ 57.5M 15.45 11.63 64.72 85.98 4.64 2.33 10242
SOFT [38] L $\mathcal{O}(Nd)$ 57.9M 12.18 5.38 82.10 185.87 3.65 1.08 10242
FLatten [21] L $\mathcal{O}(Nd^{2})$ 57.6M 14.92 10.85 67.02 92.17 4.48 2.17 10242
Agent Attn [23] L $\mathcal{O}(Nd)$ 58.1M 12.35 5.45 80.97 183.49 3.71 1.09 10242
MLLA [22] L $\mathcal{O}(Nd^{2})$ 57.5M 14.68 10.52 68.12 95.06 4.40 2.10 10242
InfViT Linear I $\bm{\mathcal{O}(N)}$ 53.5M 9.41 4.33 106.27 230.95 2.82 0.87 92162
Extreme-resolution stress test (Linear InfViT only — all others OOM)
InfViT Linear (4L) I $\mathcal{O}(N)$ 53.5M 199.22a 320.27b 5.02 3.12 59.77 64.05 92162
InfViT Linear (24L) I $\mathcal{O}(N)$ 305.4M 1783.17b 0.56 356.63 92162
Table 2 reports latency, throughput, and energy for nine attention mechanisms at $1024^{2}$ resolution. The sub-quadratic methods form three clearly separated tiers based on their asymptotic behaviour:
Tier 1: $\mathcal{O}(Nd^{2})$ feature-map attention (Performer, FLatten, MLLA). These methods replace the $N{\times}N$ softmax matrix with a product of $N{\times}d$ feature maps, but the outer-product state update introduces an $\mathcal{O}(d^{2})$ factor per token. In the 4L-64H configuration ($d_{h}{=}12$), the low head dimension makes this cost manageable (65–95 img/s), but in the 24L-16H setting ($d_{h}{=}48$) the $d^{2}$ factor grows $16{\times}$ and throughput drops to 17–25 img/s—barely faster than softmax.
Tier 2: $\mathcal{O}(Nd)$ projection-based (Linformer, SOFT, Agent Attn). By projecting to a fixed rank or aggregating through agent tokens, these methods eliminate the $d^{2}$ dependence, yielding throughput of 178–186 img/s in the 4L setting. However, the per-token cost still multiplies $d$ by the sequence length at each head, limiting the absolute advantage over Tier 1 at modest resolutions.
[…]
Extreme resolution. At $9216^{2}$ with patch size 16, the sequence length reaches N=331,776N{=}331{,}776 tokens. All methods except InfViT Linear trigger out-of-memory errors before reaching this resolution—even FlashAttention-assisted softmax ViT, which eliminates the $\mathcal{O}(N^{2})$ memory wall but retains $\mathcal{O}(N^{2})$ compute. InfViT Linear 4L sustains 3.12 img/s and 64.05 J/img at $9216^{2}$; the 24L variant reaches 0.56 img/s—slow in absolute terms, but the only model that completes the computation at all. As shown in Fig. 2(b), InfViT Linear occupies the top-left corner of the throughput-vs-energy space, clearly separated from all $\mathcal{O}(Nd)$ and $\mathcal{O}(Nd^{2})$ baselines.
4.1.1 4L vs. 24L Depth Comparison.
Increasing depth from 4 to 24 layers (Fig. 5) reduces absolute throughput for all models—InfViT Linear drops from 231 to $\sim$40 img/s—because deeper networks incur proportionally more overhead in layer norms, FFN blocks, and residual projections, which are $\mathcal{O}(Nd)$ per layer and dominate at small $N$. Crucially, the relative ranking is preserved: InfViT Linear leads in both throughput (left panel) and energy efficiency (right panel) at both depths. The $\mathcal{O}(Nd)$ baselines (Linformer, SOFT, Agent Attn) cluster at $10{-}11{\times}$ speed-up over Standard ViT, while the $\mathcal{O}(Nd^{2})$ methods (Performer, FLatten, MLLA) reach only $5{-}6{\times}$. This confirms that the complexity advantage of Linear-InfSA is not an artifact of a particular architecture size: it holds across both shallow, lightweight configurations and deep, high-capacity ones.
4.1.2 Additional Efficiency Breakdowns.
(a) Throughput
(b) Speed-up
(c) Energy per image
Figure 6: Efficiency breakdown (4L-64H, inference at $\mathbf{1024^{2}}$). (a) Absolute throughput (img/s); (b) speed-up relative to Standard ViT; (c) energy per image (J/img). InfViT Linear achieves the highest throughput and lowest energy consumption across all methods.
Fig. 6 disaggregates the 4L-64H efficiency metrics into three complementary views:
  • Throughput (Fig. 6(a)): InfViT Linear processes 231 img/s, versus 178–186 img/s for the best $\mathcal{O}(Nd)$ methods and 65–95 img/s for $\mathcal{O}(Nd^{2})$ methods. The 17 img/s of Standard ViT is $13{\times}$ slower.
  • Relative speed-up (Fig. 6(b)): Normalized to Standard ViT, InfViT Linear achieves $13.4{\times}$, while the next-fastest method (SOFT) reaches $10.8{\times}$. The $\mathcal{O}(Nd^{2})$ group (Performer, FLatten, MLLA) plateaus at $5{-}6{\times}$, illustrating the practical cost of the $d^{2}$ factor.
  • Energy per image (Fig. 6(c)): InfViT Linear consumes 0.87 J/img, roughly half the next-best methods (1.08–1.13 J for $\mathcal{O}(Nd)$ baselines). Projected over a large-scale dataset—for example, 1.28M ImageNet training images per epoch—this translates to $\sim$1.1 kJ vs. $\sim$1.4 kJ per epoch at $1024^{2}$ resolution, a saving that compounds over 300 training epochs.
The three panels together confirm that the $\mathcal{O}(N)$ theoretical complexity of Linear-InfSA translates faithfully into wall-clock and energy advantages and is not masked by constant factors or implementation overhead.
Figure 7: Training vs. inference latency (4L-64H, $\mathbf{1024^{2}}$). Paired bars show training latency (left, darker) and inference latency (right, lighter) per mechanism. Training includes the full forward–backward–update cycle in FP16 mixed precision; inference is forward-only. InfViT Linear achieves 9.41 ms (train) and 4.33 ms (infer)—the lowest in both regimes. The $2.2{\times}$ train/infer ratio is consistent with the rule of thumb that backward passes cost roughly twice the forward pass.
Fig. 7 presents a per-mechanism latency bar chart comparing training and inference times in the 4L-64H configuration. InfViT Linear achieves 9.41 ms (train) and 4.33 ms (infer), surpassing all baselines. The gap between training and inference latency reflects the cost of the backward pass and gradient accumulation; for InfViT Linear this overhead is minimal because the attention mechanism involves only vector reductions and broadcasts, which have simple and efficient gradients. For softmax-based methods, the backward pass must additionally recompute or store the $N{\times}N$ attention matrix, explaining their disproportionately higher training latency.
4.2 Attention Quality: Degradation and Localization
Figure 8: Attention quality summary. MoRF-AOC, ROC-AUC, and PR-AUC (%). Both InfSA variants outperform Standard ViT by 20–34 pp.
We evaluate whether InfSA attention maps are semantically grounded—i.e., whether high-attention patches correspond to object regions rather than diffuse background. This is important because interpretability is a core design goal of InfSA: the graph-centrality formulation should, by construction, assign high importance to the most structurally central (object) tokens. We use four complementary protocols to test this hypothesis.
[…]
MoRF degradation [51] progressively removes patches in order of
emphdecreasing attention and tracks the drop in classification confidence (Area Over the Curve, AOC $\uparrow$). If attention concentrates on discriminative regions, removing them should quickly destroy the model’s prediction. Linear InfSA achieves the steepest MoRF drop (AOC == 76.0%), followed by Pure InfSA (71.7%); Standard ViT’s nearly flat curve (AOC == 42.6%) reveals that its attention is so diffuse that removing high-attention patches barely affects confidence—a sign that softmax attention spreads weight over background rather than foreground. For LeRF (Least Relevant First), Pure InfSA maintains the highest retention (AUC == 65.3%), indicating that its
emphlow-attention patches are truly irrelevant.
Bounding-box localization treats the attention map as a patch-level binary detector and evaluates it against ImageNet ground-truth bounding boxes (2
,088 validation images with annotations). Pure InfSA leads in ROC-AUC (77.3
These improvements are large: a 20+20{+}
,pp gap in PR-AUC means that InfSA’s attention map can serve as a competitive object detector even though it was never trained with localization supervision. Both evaluations confirm that InfSA produces spatially selective, semantically aligned attention (Fig.
reffig:attn-quality).
4.2.1 Full MoRF, LeRF, ROC, and Precision–Recall Curves.
Figs. 9 and 10 provide the complete curve-level evaluations for all four attention quality protocols.

### v2

@@ 3.3 Absorbing Markov Chain Interpretation of InfSA
The Neumann-series formulation of Pure InfSA (Eq. 10) admits a direct probabilistic interpretation through absorbing Markov chains [Roffo9119168, kemeny1960finite, seneta2006nonnegative]. This connection shows that InfSA computes expected visit counts in a random walk over the token graph, linking self-attention to classical stochastic processes.
@@
where $\mathbf{M}_{ij}\geq 0$ represents the transition probability from token $i$ to token $j$, and $R_{i}\geq 0$ is the per-step absorption probability at token $i$. The matrix $\mathbf{M}$ is substochastic ($\mathbf{M}\mathbf{1}\leq\mathbf{1}$, with strict inequality for at least one row) whenever $\gamma\max_{i}\sigma_{i}<1$, where $\sigma_{i}=\sum_{j}\hat{A}_{ij}$. Frobenius normalization constrains $\|\hat{A}\|_{F}=1$, which ensures $\rho(\mathbf{M})=\gamma\,\rho(\hat{A})<1$. We augment the $N$ token states with a single absorbing state $\mathfrak{a}$, yielding the canonical form [kemeny1960finite]:
@@
which is precisely the Neumann kernel already underlying InfSA (Eq. 10). The entry $N_{ij}$ has a concrete probabilistic meaning [kemeny1960finite, seneta2006nonnegative]: it equals the expected number of times the random walk visits token $j$ before absorption, given that it starts at token $i$. The InfSA path-integral matrix $S=N-I=(I-\gamma\,\hat{A})^{-1}-I$ therefore counts expected visits excluding the starting state.
@@
These scores coincide with Katz centrality [katz1953new], confirming that InfSA ranks tokens by global structural role rather than by local query–key affinity.
Why Frobenius normalization enables absorption. Softmax normalization gives row-stochastic matrices ($\hat{A}\mathbf{1}=\mathbf{1}$), corresponding to a closed Markov chain with no absorbing state—the source of oversmoothing [li2018deeper]. Frobenius normalization breaks row-stochasticity ($\rho(\hat{A})<1$ in practice), introducing a positive absorption probability at every step and ensuring convergence. This absorbing-chain view is structurally identical to the Markov interpretation of Roffo et al. [Roffo7410835, Roffo9119168], where features are transient states and the fundamental matrix ranks them by expected walk persistence. InfSA extends this principle from feature graphs to token graphs: a token receives high centrality when it participates in many long, likely walks before diffusion terminates.
@@ 3.4 Linear-InfSA: Efficient Centrality Approximation
providing a positive signal that reflects token prominence in embedding space [romero2021geometric]. These energies are normalized without softmax:
@@
yielding a soft importance score that serves as a proxy for the dominant eigenvector of a positive operator [katz1953new, horn2012matrix]. No softmax or Frobenius normalization is required; the $\ell_{1}$ constraint ensures numerical stability.
@@
Query-independent weighting by design. The weight vector $a$ depends only on Keys (equivalently Queries, since Q=KQ{=}K) and is shared across all positions—a direct consequence of the eigenvector limit, where the operator is dominated by its principal eigenvector $v$, producing an effective rank-1 matrix $vv^{\top}$. We recover expressivity by using 64 heads (vs. 16 in standard ViT), so that each head captures a different centrality mode; on frozen checkpoints, $a$ achieves 0.985 mean cosine similarity with the true Perron eigenvector (Supplement Sec. 3). Each Linear-InfSA head produces a rank-1 output (shared context broadcast to all positions). With 64 heads and per-head dimension $d_{h}{=}12$, the concatenated output spans a $64\times 12=768$-dimensional subspace, matching standard ViT capacity.
Linear-InfSA avoids constructing the full attention matrix, reducing complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. The vector $a$ is a linear-time surrogate for the dominant eigenvector [seneta2006nonnegative, lemmens2012nonlinear], obtained classically as $v=\lim_{k\to\infty}A^{k}x_{0}/\|A^{k}x_{0}\|_{1}$; this $k{\to}\infty$ characterization motivates the term “infinite” in Linear-InfSA. The ReLU in Eq. (19) suppresses anti-aligned tokens, and under nonnegativity the map defines a positively 1-homogeneous operator whose iterates converge to a unique direction by nonlinear Perron–Frobenius theory [lemmens2012nonlinear, birkhoff1957extensions] (see Supplement for the full convergence argument).
@@
Relation to global pooling. Although the per-head output is a broadcast context vector, Linear-InfSA is not reducible to generic global pooling [hu2018squeeze] or additive-attention aggregation [wu2021fastformer]: the weight vector $a$ is the unique fixed point of a nonlinear Perron–Frobenius operator whose spectrum encodes the full inter-token similarity structure, not a learned or heuristic saliency score. This spectral grounding is verified empirically—$a$ achieves 0.985 mean cosine similarity with the true Perron eigenvector (Supplement Sec. 3)—and ensures that the aggregation captures global graph centrality rather than local token prominence.
4 Experiments
We evaluate InfSA by integrating it into a 4-layer Vision Transformer (ViT) with patch size 16. Pure InfViT uses 16 full-attention heads, while Linear InfViT uses 64 lightweight heads. Both models have comparable compute and parameter counts (58M vs. 53M). Full training details are in the Supplement.
4.1 Scalability to Extreme Input Resolutions
[…]
Train Infer Train Infer Train Infer
24-layer, 16-head configuration ($d_{h}{=}48$) — fair depth comparison
Linformer [linformer] L $\mathcal{O}(Nd)$ 331.5M 55.64 38.95 17.97 25.67 16.69 7.79 10242
Performer [performer] L $\mathcal{O}(Nd^{2})$ 330.3M 58.14 42.80 17.20 23.36 17.44 8.56 10242
SOFT [lu2021soft] L $\mathcal{O}(Nd)$ 330.8M 56.10 36.44 17.83 27.44 16.83 7.29 10242
FLatten [han2023flatten] L $\mathcal{O}(Nd^{2})$ 330.5M 57.82 40.15 17.29 24.91 17.35 8.03 10242
Agent Attn [han2024agent] L $\mathcal{O}(Nd)$ 331.0M 56.55 35.81 17.68 27.93 16.97 7.16 10242
MLLA [han2024mlla] L $\mathcal{O}(Nd^{2})$ 330.4M 57.38 39.52 17.43 25.30 17.21 7.90 10242
[…]
4-layer, 64-head configuration ($d_{h}{=}12$) — proposed lightweight
Linformer [linformer] L $\mathcal{O}(Nd)$ 58.3M 12.49 5.63 80.04 177.62 3.75 1.13 10242
Performer [performer] L $\mathcal{O}(Nd^{2})$ 57.5M 15.45 11.63 64.72 85.98 4.64 2.33 10242
SOFT [lu2021soft] L $\mathcal{O}(Nd)$ 57.9M 12.18 5.38 82.10 185.87 3.65 1.08 10242
FLatten [han2023flatten] L $\mathcal{O}(Nd^{2})$ 57.6M 14.92 10.85 67.02 92.17 4.48 2.17 10242
Agent Attn [han2024agent] L $\mathcal{O}(Nd)$ 58.1M 12.35 5.45 80.97 183.49 3.71 1.09 10242
MLLA [han2024mlla] L $\mathcal{O}(Nd^{2})$ 57.5M 14.68 10.52 68.12 95.06 4.40 2.10 10242
[…]
Extreme-resolution stress test (Linear InfViT only — all others OOM)
[…]
4.2 Attention Quality: Degradation and Localization
Figure 5: Attention quality summary. MoRF-AOC, ROC-AUC, and PR-AUC (%). Both InfSA variants outperform Standard ViT by 20–34 pp. Full curves in the Supplement (Figs. 78).
We evaluate whether InfSA attention maps are semantically grounded on the ImageNet-1K validation set. Attention maps are extracted from 24-layer models with each respective attention mechanism (ViT-L/16 backbone, $14{\times}14$ tokens, 304M params), following standard practice for interpretability evaluation.
MoRF degradation [samek2016evaluating] progressively removes high-attention patches and tracks confidence drop (Area Over the Curve, AOC $\uparrow$). Linear InfSA achieves the steepest MoRF drop (AOC == 76.0%), followed by Pure InfSA (71.7%); Standard ViT’s flat curve (AOC == 42.6%) reveals diffuse attention. For LeRF (Least Relevant First), Pure InfSA maintains the highest retention (AUC == 65.3%).
Bounding-box localization treats attention as a patch-level detector against ImageNet ground-truth boxes (2 088 images). Pure InfSA leads in ROC-AUC (77.3%) and Linear InfSA in PR-AUC (76.1%), versus 53.8% and 56.2% for Standard ViT (++20–24 pp). Both evaluations confirm that InfSA produces spatially selective, semantically aligned attention (Fig. 5).
4.3 ImageNet Classification Results
All models are trained on ImageNet-1K (1.28M images, 300 epochs, batch 64) with a DeiT-style recipe [touvron2021training], without external data, distillation, or self-supervised pretraining. Our ViT baseline reaches 81.5%; the +3.2+3.2 pp gain of Linear InfViT 4L (84.7%, 53.5M params, 59 GFLOPs) is purely architectural. Note that InfViT-4L operates at 59 GFLOPs ($224^{2}$) vs. 17.5 GFLOPs for ViT-B/16, owing to its 64-head design. However, the $\mathcal{O}(N)$ complexity of Linear-InfSA means this gap narrows at higher resolutions and reverses beyond $\sim$$512^{2}$, where quadratic baselines’ FLOPs grow as $N^{2}$ while InfViT’s grow linearly. Fig. 6 plots ImageNet-1K and ImageNet-V2 [recht2019imagenet] top-1 accuracy against parameter count. On IN-1K, InfViT Pure 24L reaches 85.4%, within 0.4 pp of RAVLT-L [fan2024ravlt] (85.8%). The 4L variants (85.1% Pure, 84.7% Linear) surpass Agent Attn-B (84.1%) and FLatten ($\leq$}84.5%) at roughly half the parameters. On IN-V2, all four InfViT models exceed every baseline (up to 79.8% vs. 76.8% for RAVLT-L), indicating robust generalization. The 24L models improve by only 0.3–0.4 pp over 4L at $6{\times}$ the parameter cost; the 4L-64H variants are the recommended configuration. The full per-method table is in the Supplement (Table 3).
Ablation results (path-decay $\gamma$ and activation function) are reported in the Supplement (Table 4). In brief: accuracy and MoRF-AOC peak at $\gamma{=}0.7$ with ReLU (84.7%, 76.0%); the Linear-InfSA weight vector achieves 0.985 cosine similarity with the Perron eigenvector of the full operator on small token sets (Supplement Sec. 3).
5 Conclusion
Empirically, InfViT models show strong performance across classification, localization, and scalability tasks. Pure InfViT improves ImageNet accuracy and attention alignment, while Linear InfViT scales to $9216^{2}$ resolution with $13{\times}$ better energy efficiency than standard ViT of equal depth. Compact 4-layer variants match or exceed larger baselines in accuracy at a fraction of the parameters, enabling practical deployment. The graph-theoretic principles underlying InfSA are modality-agnostic, suggesting natural extensions to NLP, multi-modal models, video understanding, and dense prediction tasks; all code and models will be released upon acceptance for reproducibility.
Acknowledgements
This work builds in part on conceptual directions previously explored in the MVL / Toyota Motor Europe collaboration.
Supplementary Material
6 Why Pure InfSA and Linear-InfSA Are “Infinite”
@@
7 Convergence of Linear-InfSA
@@
By the nonlinear Perron–Frobenius theorem for positively 1-homogeneous, order-preserving maps on the positive cone [lemmens2012nonlinear, birkhoff1957extensions], the normalized iterates $F^{k}(x_{0})/\|F^{k}(x_{0})\|_{1}$ converge to a unique eigenvector direction. In the strictly nonnegative case (ensured by $\varepsilon>0$), this reduces exactly to the classical power method, and the fixed point is the Perron eigenvector of the implicit positive operator. In all experiments we use $\varepsilon=10^{-6}$. Our empirical validation (Sect. 8) confirms that the single-step Linear-InfSA weight vector $a$ achieves 0.985 mean cosine similarity with this eigenvector.
8 Eigenvector Alignment Validation
Experimental setup.
Ground-truth Perron eigenvector.
Reconstruction of Linear-InfSA weights.
From the same tensors, we reconstruct the weight vector $a$ using the closed-form equations from the main paper:
$a_{t}$ $=\frac{s_{t}}{\textstyle\sum_{s}s_{s}+\varepsilon}.$ (27)
Metrics.
Metric Mean Std. dev.
Cosine similarity $0.985$ $0.036$
Spearman correlation $0.937$ $0.132$
Discussion.
Table 2 shows that the Linear-InfSA weights are almost perfectly aligned with the Perron eigenvector: mean cosine 0.985 and mean Spearman 0.937. This validates that Linear-InfSA faithfully preserves the dominant spectral structure of the full operator, justifying its use as a drop-in $\mathcal{O}(N)$ replacement.
9 Additional Attention Quality Curves
The main paper (Sect. 4.2, Fig. 5) presents a compact bar-chart summary of attention quality. Here we provide the full curve-level evaluations: MoRF degradation, ROC, LeRF retention, and Precision–Recall (PR), all evaluated under the same protocol.